OpenStudy (cutiecomittee123):

log 55 is the power to which ______ must be raised inorder to produce a value of 55.

jimthompson5910 (jim_thompson5910):

If the base isn't stated, then by default the log is base 10 ie \[\Large \log(55) = \log_{10}(55)\]

jimthompson5910 (jim_thompson5910):

the base of the log will also be the base of the exponential expression

OpenStudy (cutiecomittee123):

Okay thats what I thought but I had to make sure. Thanks:) If not Im thinking it is just "x". 10 sounds like a much more accurate answer.

jimthompson5910 (jim_thompson5910):

it's 10 because you would do \[\Large 10^{\log(55)}\] which gives 55 as the answer

jimthompson5910 (jim_thompson5910):

or put another way, log(55) answers the question: "what is the value of x in \(\Large 10^x = 55\) ?"